Hi-Tech Mortgage Company uses a process costing system to accumulate costs in its loan application department. When an application is completed, it is forwarded to the loan department for final processing. The following processing and cost data pertain to September:

1. Applications in process on Beginning work in process:
September 1, 80      Direct materials $1,160
2. Applications started in      Conversion costs 3,820
September, 1,110 September costs:
3. Completed applications during      Direct materials $4,790
September, 860      Direct labour 11,800
4. Applications still in process at September 30 were      Overhead 14,004
100% complete in terms of materials (forms) and 60%
complete in terms of conversion costs.


Materials are the forms used in the application process, and these costs are incurred at the beginning of the process. Conversion costs are incurred uniformly during the process.

Determine the equivalent units of service (production) for materials and conversion costs.

Materials Conversion
Costs
Total equivalent units
Answer #1

Units in ending inventory = Beginning units + Started units - Completed units

= 80 + 1,110 - 860

= 330

Equivalent units = Units completed + (Completed units X Percentage of completion)

Materials = 860 + (330 X 100%) = 1,190

Conversion costs = 860 + (330 X 60%) = 1,058